# Enhancing LLM Reliability via Explicit Knowledge Boundary Modeling

### v2

@@ Abstract
Large language models (LLMs) frequently hallucinate due to misaligned self-awareness, generating erroneous outputs when addressing queries beyond their knowledge boundaries. While existing approaches mitigate hallucinations via uncertainty estimation or query rejection, they suffer from computational inefficiency or sacrificed helpfulness. To address these issues, we propose the Explicit Knowledge Boundary Modeling (EKBM) framework, integrating fast and slow reasoning systems to harmonize reliability and usability. The framework first employs a fast-thinking model to generate confidence-labeled responses, enabling immediate use of high-confidence outputs. For uncertain predictions, a slow refinement model conducts targeted reasoning to improve accuracy. To align model behavior with our proposed object, we propose a hybrid training pipeline, enhancing self-awareness without degrading task performance. Evaluations on dialogue state tracking tasks demonstrate that EKBM achieves superior model reliability over uncertainty-based baselines. Further analysis reveals that refinement substantially boosts accuracy while maintaining low computational overhead. Our work establishes a scalable paradigm for advancing LLM reliability and balancing accuracy and practical utility in error-sensitive applications.
Enhancing LLM Reliability via Explicit Knowledge Boundary Modeling
Hang Zheng1††thanks: Equal contribution., Hongshen Xu1, Yuncong Liu1, Lu Chen1, Pascale Fung2, Kai Yu1 1X-LANCE Lab, Department of Computer Science and Engineering, Shanghai Jiao Tong University 2Center for Artificial Intelligence Research (CAiRE), Hong Kong University of Science and Technology {azure123, xuhongshen, chenlusz, kai.yu}@sjtu.edu.cn
@@ 1 Introduction
Recently, large language models (LLMs) have exhibited impressive text generation capabilities Abdullah et al. (2022). However, despite their advancements, LLMs are prone to hallucination, where generated content misaligns with context or factual information Zhang et al. (2023). Such inaccuracies can be particularly harmful in low error-tolerance applications, eroding user trust in LLM reliability.
@@
Extensive works have been devoted to mitigating hallucinations in LLMs Tonmoy et al. (2024). Hallucinations often arise from misalignments between an LLM’s outputs and its intrinsic knowledge boundaries, leading to errors when models attempt to answer questions beyond their expertiseLi et al. (2024b). Improving the model’s self-awareness—its capacity to accurately assess its own knowledge boundaries and outputs correctness, can effectively mitigate this misalignment. Several uncertainty-based methods have been proposed to estimate model confidence and indirectly reflect the model’s self-awareness Huang et al. (2024). However, these approaches are not inherent to the model’s endogenous capabilities and frequently suffer from high computational costs or instability across thresholds. Alternative strategies emphasize aligning knowledge boundaries with actions by rejection of out-of-scope queries Xu et al. (2024b), avoiding mistakes but sacrificing helpfulness—a trade-off that undermines model utility particularly in complex, multi-step tasks where partial correctness remains valuable.
@@
To mitigate these limitations, we propose a novel alignment objective: improving model awareness for knowledge boundaries, enabling models to respond maximally while explicitly distinguishing outputs of high and low confidence. We introduce the concepts of "sure" and "unsure" to clarify this distinction. A reliable system should deliver near-perfect accuracy for "sure" predictions and provide helpful information in the “unsure” category, serving as a form of soft rejection, appropriately managing user expectations regarding accuracy.
@@
Building on this concept, we propose the Explicit Knowledge Boundary Modeling (EKBM) framework. EKBM integrates both Fast and Slow systems: a fast-thinking model generates responses annotated with confidence labels, allowing immediate utilization of high-confidence outputs, while a slow-thinking refinement model engages in deliberate reasoning to enhance accuracy of low-confidence predictions. Crucially, EKBM does not merely filter uncertain outputs but leverages them as opportunities for improvement, ensuring both precision and coverage. The collaboration between these two systems strikes a balance between reliability and efficiency, with the self-awareness capability of the fast-thinking model being essential for the system’s effective operation.
@@
To operationalize this paradigm, we design a training pipeline that aligns the model’s capability boundaries and improves self-awareness. Evaluations on dialogue state tracking (DST) tasks demonstrate that our method effectively enables the model to reliably and accurately categorize its outputs into "sure" and "unsure" categories compared to baseline methods. Additionally, we develop an automatic refinement model for "unsure" outputs, showing that the EKBM system achieve superior performance compared to traditional supervised fine-tuning (SFT) approaches. Our analysis of various training methods and model behaviors provides further valuable insights.
@@
2 Related Work
The knowledge boundary of LLMs has emerged as a critical topic in recent research, highlighting their limitations in generating reliable outputs Yin et al. (2024); Li et al. (2024b). Misalignment between model behavior and knowledge boundaries can result in factual hallucinations Huang et al. (2023) and ambiguous responses Liu et al. (2023).
Various methods have been proposed to assess these boundaries. Uncertainty-based approaches quantify prediction confidence through token probabilities Manakul et al. (2023) and consistency Chen and Mueller (2024). Calibration strategies, including prompting Tian et al. (2023) and fine-tuning Tao et al. (2024) for improved confidence expression, align model confidence with prediction accuracy. Internal State Probing evaluates prediction factuality by analyzing model states, such as activations Li et al. (2024a). While effective, these methods often rely on external post-hoc techniques, which are computationally expensive and lack integration with the model’s endogenous reasoning processes Huang et al. (2024); Zhou et al. (2024b).
To address this misalignment, researchers have implemented knowledge-enhanced fine-tuning Zhou et al. (2024a) and retrieval-augmented generation (RAG, Lewis et al., 2020) to bolster model capabilities, thereby reducing hallucinations. Some approaches also guide models in self-reflection to minimize self-contradictions Wei et al. (2022); Ji et al. (2023), albeit at the cost of increased overhead. For queries beyond the model’s capabilities, refusal or rejection strategies can prevent misinformation Xu et al. (2024b); Chen et al. (2024), though they may inadvertently reduce overall helpfulness.
@@ 3.1 Alignment for Reliability
Recent studies have introduced rejection as a mechanism to help the model differentiate between answerable (in-boundary) and unanswerable (out-of-boundary) queries. For unanswerable queries, the model is encouraged to refrain from responding. The corresponding alignment objective can be articulated as:
@@
While this rejection strategy aligns with reliability goals, it may unduly limit the model’s helpfulness in practical scenarios. Therefore, we propose a more nuanced objective: the model should aim to provide answers whenever feasible, categorizing outputs into high-confidence ("sure") and low-confidence ("unsure") responses. The model must ensure high accuracy in the "sure" category, while tolerating some errors in the "unsure" category, which can serve as references for users or inputs for further refinement. Our proposed alignment objective is formalized as:
Here, and represent correct and incorrect responses, respectively, while and represent "sure" and "unsure" as levels of confidence.
This objective guarantees that correct "sure" predictions are prioritized, followed by correct "unsure" predictions, which, although less confident, retain utility. Incorrect predictions in the "unsure" category are discouraged but more acceptable than those in the "sure" category, as the latter significantly undermine reliability.
@@ 3.2.1 Weighted-F1
We modify the F1 score to adapt to the alignment objective of our EKBM framework. Since predictions labeled as "unsure" represent an intermediate state of uncertainty, we assign a weighted contribution (denoted by and ) to predictions in this category. The weights reflect the partial utility of "unsure" predictions as potential candidates for refinement. The modified precision and recall are defined as:
Where STP and SFP represent Sure True Positives and Sure False Positives, while UTP and UFP denote the corresponding cases of Unsure. For UTP in Recall calculation, we only require that a prediction is partially correct (able to be refined). The modified Weighted-F1 score is defined as:
The design allows flexible handling of "unsure" predictions, effectively managing their intermediate utility. influences the calculation of Precision, reflecting the model’s tolerance for errors among "unsure" predictions; as increases, tolerance for such errors decreases. Conversely, affects Recall, indicating the degree of reward for providing additional reference information for "unsure" predictions. As increases, the model is encouraged to enhance helpfulness by generating more "unsure" predictions as a basis for refinement.
We discuss a few corner cases: (1) = 0, = 0: Weighted-F1 disregards all "unsure" predictions. (2) = 0, = 1: Errors in the "unsure" category are ignored, rewarding effective recall, representing the theoretical upper limit post-perfect refinement. (3) = 1, = 0: Lacks significant interpretation. (4) = 1, = 1: Eliminates confidence labeling, reverting to traditional methods.
3.3 Reliability Metrics
We employ two specific Weighted-F1 values as performance metrics:
Optimal-F1 Defined as , representing the theoretical upper limit after perfect refinement of "unsure" predictions.
Quality-F1 Defined as , it assigns "unsure" predictions half the weight of "sure" ones, encouraging the model to select "sure" when confident while minimizing penalties for "unsure" outputs. Quality-F1 measures the model’s self-awareness in accurately classifying both types.
Among these, Quality-F1 is the core metric for assessing model reliability, while Optimal-F1 serving as a supplementary reference.
@@ 4.1 Explicit Knowledge Boundary Modeling
We formalize our proposed framework, Explicit Knowledge Boundary Modeling (EKBM), which aims to improve the reliability and self-awareness of the model by explicitly modeling its knowledge boundaries and improve overall performance by incorporating a refinement model to improve the "unsure" predictions. As shown in Figure 2 (1), the framework divides the model’s decision-making process into two distinct stages.
@@
In the first stage, the model performs the primary task while simultaneously assigning a confidence label ("sure" or "unsure") to each prediction. Predictions labeled as "sure" are directly accepted as final outputs, ensuring high precision.
@@
In the second stage, the "unsure" predictions are further refined. This refinement process can involve various strategies, such as user confirmation, automated multi-step reasoning, or other post-processing techniques. This additional stage incurs extra cost but increases the overall accuracy and reliability of the system.
Our two-stage framework balances immediate usability with comprehensive coverage, making the model both reliable and helpful.
@@ 4.2 Reliability Alignment
To align the model’s behavior with our proposed alignment objective, we explore various training approaches to enhance its intrinsic self-awareness Xu et al. (2025). To minimize interference from task performance improvements, we introduce minimal supervisory signals related to task knowledge. We employ Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO), with the data construction process illustrated in Figure 2 (2).
@@
We assume that task-specific knowledge boundary of the model can be approximated by repeatedly sampling the model’s outputs on the training data. Based on this, our SFT data construction approach is intuitive. We conduct multiple rounds of sampling for each training sample. Predictions that are consistently accurate are labeled as "sure," while those exhibiting errors or omissions receive an "unsure" label. Importantly, we refrain from correcting erroneous outputs to avoid introducing supervisory signals that could compromise the model’s task performance.
The number of sampling rounds is a hyperparameter denoted as i. A larger i results in a more conservative model, as the proportion of "unsure" labels in the dataset increases with greater i. In our study, we set i = 1 by default to accurately reflect the model’s realistic capability boundary.
@@
DPO enables the model to learn directly from preference pairs, making the design of an effective preference strategy crucial. Given our alignment objective, it is intuitive to use Weighted-F1 as the preference score function. The values in Weighted-F1 significantly influence the model’s behavior and performance. Through experimentation (see Section 5.6), we determined that setting to 0.25 and to 0.75 best balances the model’s immediate helpfulness and its potential performance after refinement. Therefore, this setting will be used by default unless specified. To enhance the stability of DPO training, we ensure that all data is generated through multi-sampling, aligning with the model’s output distribution.
@@ 4.3 Refinement Model
We train automated refinement models for each dataset to further optimize by reasoning on unsure predictions, utilizing supervised fine-tuning (SFT) based on the LLAMA3 8B model. Overall, our refinement process operates at a fine-grained level, with each unsure prediction being individually refined. We employ a multi-step reasoning paradigm similar to the Chain of Thought Wei et al. (2022) methodology and use GPT-4o Hurst et al. (2024) to generate the reasoning process. This process is illustrated in Figure 2 (3), and detailed information can be found in Appendix A.
@@ 5 Experiments
RQ1 Does our proposed reliability training pipeline outperform current methods in enhancing self-awareness abilities?
@@
RQ3 Does our approach demonstrate strong scalability, performing well across foundation models with various performance?
@@ 5.1.1 Model and Baselines
For our experiments, we utilize the LLaMA3 8B model and the Qwen2.5 7B model as the backbone.
For reliability alignment, we conducted Reliability-SFT and integrated DPO to create two additional variants: Reliability-SFT+DPO Joint Training, which undergoes joint training with SFT and DPO, and Reliability-SFT+DPO Post Training, where DPO training follows initial SFT training, optimizing the model beyond the Reliability-SFT foundation. The key distinction lies in the different source distributions of preference data sampled for DPO.
To evaluate the reliability of our proposed methods, we compare their performance against several baseline approaches, including three prompt-based and three uncertainty-based methods.
@@
methods include three approaches: Direct, Verbose, and Self-Reflection (SR). The Direct method predicts outputs without incorporating any confidence labels, while Verbose generates confidence labels alongside predictions based on predefined principles. SR evaluates prediction reliability through individual reflections.
@@
methods consist of three approaches: Token Probability (Prob), Self-Consistency (SC), and P(True) Kadavath et al. (2022). These methods classify outputs as "sure" or "unsure" using different techniques. Prob calculates average token-level probabilities, SC assesses the frequency of predictions across multiple samples, and P(True) determines the probability of the "True" token during model self-evaluation.
The details of the baselines can be found in E.1.

### v3

@@ Abstract
Large language models (LLMs) are prone to hallucination stemming from misaligned self-awareness, particularly when processing queries exceeding their knowledge boundaries. While existing mitigation strategies employ uncertainty estimation or query rejection mechanisms, they suffer from computational efficiency and sacrificed helpfulness. To address these issues, we propose the Explicit Knowledge Boundary Modeling (EKBM) framework, integrating fast and slow reasoning systems to harmonize reliability and usability. The framework first employs a fast-thinking model to generate confidence-labeled responses, enabling immediate utilization of high-confidence outputs, whereas uncertain predictions trigger a slow refinement model for accuracy improvement. To align model behavior with our proposed object, we propose a hybrid training pipeline, enhancing self-awareness without degrading task performance. Evaluations on dialogue state tracking tasks demonstrate that EKBM achieves superior model reliability over uncertainty-based baselines. Further analysis reveals that refinement substantially boosts accuracy while maintaining low computational overhead. The framework establishes a scalable paradigm for deploying reliable LLMs in error-sensitive applications, effectively balancing accuracy and practical utility.
@@ 1 Introduction
Recently, large language models (LLMs) have demonstrated impressive text generation capabilities (Abdullah et al., 2022). However, LLMs are susceptible to hallucinations, where generated content misaligns with context or factual information (Zhang et al., 2023). Hallucinations can be particularly detrimental in applications with low tolerance for error, eroding user trust in LLM reliability.
@@
Extensive works have been devoted to mitigating hallucinations in LLMs (Tonmoy et al., 2024). Hallucinations often arise from misalignments between an LLM’s outputs and its intrinsic knowledge boundaries, particularly when when addressing queries beyond its expertise (Li et al., 2024b). Improving the model’s self-awareness—its capacity to accurately assess its own knowledge boundaries and outputs correctness, can effectively mitigate this misalignment. Uncertainty-based methods estimate model confidence and indirectly reflect the model’s self-awareness (Huang et al., 2024), facing challenges of inherence to the model’s endogenous capabilities, frequently sufferring from high computational costs or instability across thresholds. Alternative strategies emphasize aligning knowledge boundaries with actions by rejection of out-of-scope queries (Xu et al., 2024b), avoiding mistakes but sacrificing helpfulness—a trade-off that undermines model utility particularly in complex, multi-step tasks where partial correctness remains valuable.
@@
To address these limitations, we propose a novel alignment objective that enhances model awareness of its knowledge boundaries, enabling models to respond maximally while explicitly distinguishing high/low confidence outputs. We introduce the concepts of ”sure” and ”unsure” to clarify this distinction. A reliable system should deliver near-perfect accuracy for ”sure” predictions and provide helpful information in the “unsure” category, serving as a form of soft rejection, appropriately managing user expectations regarding accuracy.
@@
Building on this concept, we propose the Explicit Knowledge Boundary Modeling (EKBM) framework. EKBM integrates both Fast and Slow systems: a fast-thinking model generates responses annotated with confidence labels, allowing for immediate use of high-confidence outputs, while a slow-thinking refinement model engages in deliberate reasoning to improve low-confidence predictions.
Crucially, EKBM does not merely filter uncertain outputs but leverages them as opportunities for improvement, ensuring both precision and coverage. The collaboration between these two systems strikes a balance between reliability and efficiency, with the self-awareness capability of the fast-thinking model being essential for the system’s effective functioning.
@@
To operationalize this paradigm, we design a training pipeline that aligns the model’s capability boundaries and improves self-awareness. Evaluations on dialogue state tracking (DST) tasks demonstrate that our method effectively enables the model to reliably and accurately distinguish the confidence levels of its output compared to baseline methods. Additionally, we develop an automatic refinement model for ”unsure” outputs. Experiments show that the EKBM system achieves superior overall performance compared to traditional approaches that directly tune models for the task.
@@ 2.1 Alignment for Reliability
Recent studies have introduced rejection as a mechanism to help the model differentiate between answerable (in-boundary) and unanswerable (out-of-boundary) queries, where model is encouraged to directly reject the unanswerable ones. The corresponding alignment objective can be articulated as:
@@
While this rejection strategy aligns with reliability goals, it may unduly limit the model’s helpfulness in practical scenarios. Therefore, we propose a more nuanced objective: the model should aim to provide answers whenever feasible, categorizing outputs into high-confidence (”sure”) and low-confidence (”unsure”) responses. The model must ensure high accuracy in the ”sure” category, while tolerating some errors in the ”unsure” category, which can serve as references for users or inputs for further refinement. Our proposed alignment objective is formalized as:
Here, and represent correct and incorrect responses, respectively, while and represent ”sure” and ”unsure” as levels of confidence. This objective guarantees that correct ”sure” predictions are prioritized, followed by correct ”unsure” predictions, which, while less confident, still provide utility. Incorrect predictions in the ”unsure” category are discouraged but deemed more acceptable than those in the ”sure” category, as the latter significantly undermine overall reliability.
@@ 2.2.1 Weighted-F1
We modify the F1 score to align with the objectives of our EKBM framework. Since ”unsure” predictions represent an state of uncertainty, we assign a weighted contribution (denoted by and , both ranging from 0 to 1) to reflect their partial utility as candidates for refinement. The modified precision and recall are defined as:
where STP and SFP denote Sure True Positives and Sure False Positives, while UTP and UFP represent the corresponding cases in the Unsure category. For UTP in the recall calculation, we only require that a prediction be partially correct (i.e., refinable). Unlike precision, we do not introduce the weighting parameter in the denominator of recall. This is because recall measures the proportion of correct predictions relative to the total number of gold labels, which remains unchanged regardless of the introduction of . Notably, incorporating does not affect the range of precision and recall, both of which remain within [0,1]. The modified Weighted-F1 score is defined as:
2.2.2 Metric-Objective Consistency: Weighted-F1 as a Parametric Alignment Measure
The Weighted-F1 metric structurally aligns with our four-tiered alignment objective (Equation (3)) through a parametric formulation that encodes hierarchical priorities. Its design ensures: (1) Strict enforcement for sure predictions: Unweighted sure terms prioritize high-confidence correctness (STP) while imposing maximal penalties for sure errors (SFP); (2) Controlled utility for unsure predictions: -scaled unsure terms modulate low-confidence outputs—attenuated UFP penalties tolerate unsure errors, whereas partial UTP rewards incentivize refinable correct predictions.
This dual-parameter mechanism enables scenario-specific adaptation: Weighted-Precision regulates error tolerance via (higher tightens unsure-error constraints) and Weighted-Recall controls reference-value incentives via (higher promotes helpfulness through unsure-correct outputs). The composite metric thus operationalizes the reliability-helpfulness trade-off, permitting calibration from conservative (, ) to exploratory (, ) strategies.
We discuss a few corner cases: (1) = 0, = 0: Weighted-F1 disregards all ”unsure” predictions. (2) = 0, = 1: Errors in the ”unsure” category are ignored, rewarding effective recall, representing the theoretical upper limit after perfect refinement. (3) = 1, = 0: Lacks significant interpretation. (4) = 1, = 1: Eliminates confidence labeling, reverting to traditional methods.
2.2.3 Reliability Metrics
We employ two specific Weighted-F1 scores as performance metrics. Among these, Quality-F1 is the core metric for assessing model reliability, while Optimal-F1 serving as a supplementary reference.
Quality-F1 Defined as , it assigns ”unsure” predictions half the weight of ”sure” ones, encouraging the model to select ”sure” when confident while minimizing penalties for ”unsure” outputs. Quality-F1 measures the model’s self-awareness ability and reliability.
Optimal-F1 Defined as , it represents the theoretical upper limit achievable after perfect refinement of ”unsure” predictions, reflecting the overall potential of the two-stage system.
@@ 3.1 Explicit Knowledge Boundary Modeling
We formalize our proposed framework, Explicit Knowledge Boundary Modeling (EKBM), which aims to improve the reliability and self-awareness of the model by explicitly modeling its knowledge boundaries and improve overall performance by incorporating a refinement model to improve the ”unsure” predictions. As shown in Figure 2 (1), the framework divides the model’s decision-making process into two distinct stages, balancing immediate usability with comprehensive coverage, making the system both reliable and helpful.
@@
In the first stage, the model performs the primary task while simultaneously assigning a confidence label (”sure” or ”unsure”) to each prediction. Predictions labeled as ”sure” are directly accepted as final outputs, ensuring high precision.
@@
In the second stage, ”unsure” predictions undergo further refinement through strategies like user confirmation, automated multi-step reasoning, or other post-processing techniques, incuring extra cost but increases overall accuracy and reliability of the system.
@@ 3.2 Reliability Alignment
To align the model’s behavior with our proposed alignment objective, we explore various training strategies to enhance its intrinsic self-awareness (Xu et al., 2025). We apply Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) for alignment, with the data construction process illustrated in Figure 2(2) and two detailed examples provided in Appendix §B.
@@
We assume that the model’s task-specific knowledge boundary can be approximated by repeatedly sampling on the training data. Accordingly, our SFT data construction approach is straightforward: we conduct multiple sampling for each training sample. Predictions consistently deemed accurate are labeled ”sure,” while those with errors or omissions are marked ”unsure.” Crucially, we intentionally refrain from correcting erroneous outputs to avoid introducing supervisory signals that might inadvertently improve the model’s task performance. That is, when the model produces an incorrect prediction, we simply assign a confidence label of ”unsure” without rectifying the error. The erroneous output remains in the final training data, reinforcing the model’s ability to recognize and categorize uncertain information rather than improving its accuracy directly.
The number of sampling rounds, denoted as i, is a hyperparameter. A larger i leads to a more conservative model, as the proportion of ”unsure” labels increases with greater i in the training corpus. In our study, we set i = 1 by default to accurately reflect the model’s realistic capability boundary.
@@
DPO enables the model to learn directly from preference pairs, making the design of an effective preference strategy crucial. Given our alignment objective, it is intuitive to use Weighted-F1 as the preference score function. The values in Weighted-F1 significantly influence the model’s behavior and performance. Through experimentation (see §4.5), we determined that setting to 0.25 and to 0.75 best balances the model’s immediate helpfulness and its potential performance after refinement. Therefore, this setting will be used by default unless specified. To enhance the stability of DPO training, we ensure that all data is generated through multi-sampling from the model itself, aligning with the model’s output distribution.
@@ 3.3 Refinement Model
We train automated refinement models for each dataset to further optimize unsure predictions by multi-step reasoning, leveraging supervised fine-tuning (SFT) based on the LLAMA3 8B model. Our refinement operates at a fine-grained level, addressing each unsure prediction individually. We employ a reasoning paradigm similar to the Chain of Thought (Wei et al., 2022) methodology and use GPT-4o (Hurst et al., 2024) to generate detailed reasoning process for each training prediction. This approach is illustrated in Figure 2 (3), and detailed information can be found in Appendix §C.
@@ 4 Experiments
RQ1 Does our reliability training pipeline outperform current methods in enhancing self-awareness?
@@
RQ3 Does our approach scale robustly across foundation models with varying capabilities?
@@ 4.1.1 Model and Baselines
We utilize the LLaMA3 8B model and the Qwen2.5 7B model as the backbone for our experiments.
For reliability alignment, we conducted Reliability-SFT and integrated DPO to create two additional variants: Reliability-SFT+DPO Joint Training, which undergoes joint training with SFT and DPO, and Reliability-SFT+DPO Post Training, where DPO training follows initial SFT training, optimizing the model beyond the Reliability-SFT foundation. The key distinction lies in the different source distributions of preference data sampled for DPO. Training details could be found in Appendix §A.
To evaluate the reliability of our proposed models, we compare them against several baselines, including three prompt-based and three uncertainty-based approaches (see Appendix §D). Notably, our focus is on the model’s self-awareness rather than task performance, so we do not include SOTA DST models for comparison. By default, we utilize only 1,000 samples for Reliability-SFT training and 2,000 for DPO training—significantly less than traditional models.
@@
methods encompass three approaches: Direct, Verbose, and Self-Reflection (SR). The Direct approach generates predictions without confidence estimation, whereas Verbose simultaneously produces confidence labels. SR assesses each prediction through self-evaluative reflection.
@@
methods include Token Probability (Prob), Self-Consistency (SC), and P(True) (Kadavath et al., 2022), each employing distinct techniques to classify outputs as ”sure” or ”unsure”. Prob calculates average token probabilities, SC measures the frequency of predictions across multiple samples, and P(True) determines the probability of the ”True” token during model self-evaluation.